OpenStudy (anonymous):

Find f (g(x)) if f(x) = 2x - 1 and g(x) = x + 2.

OpenStudy (foolaroundmath):

f(g(x)) = f(x+2) = 2(x+2) - 1 f(g(x)) = 2x + 3

Parth (parthkohli):

First of all - \( \color{Black}{\Rightarrow f(g(x)) = f(x + 2) }\) now \( \color{Black}{\Rightarrow f(x + 2) = 2(x + 2) - 1}\) Can you do it?